OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 5x + y = 5 3x + 2y = 3

OpenStudy (anonymous):

(1,0)

OpenStudy (anonymous):

y=5-5x 3x+10-10x=3 -7x=-7 x=1 y=5-5*1 Y=0